# For referencing this work, please cite

### v2

@@ Abstract
unsupervised, highlight detection, multimodal, representation activation sequence.
@@ III-E The Auxiliary Task of Masked FVS Reconstruction
GIFs
LSVM
RARE
LM
MINI-Net
LR
CHD
Ours
dog
0.308
0.60
0.49
0.579
0.577
0.554
0.606
0.642
gymnast.
0.335
0.41
0.35
0.417
0.574
0.623
0.711
0.758
parkour
0.540
0.61
0.50
0.670
0.698
0.701
0.742
0.709
skating
0.554
0.62
0.25
0.578
0.522
0.691
0.498
0.456
skiing
0.328
0.36
0.22
0.486
0.539
0.601
0.682
0.665
surfing
0.541
0.51
0.49
0.651
0.593
0.598
0.685
0.667
Average
0.464
0.54
0.38
0.564
0.584
0.630
0.654 0.651
@@ IV-A Evaluation datasets
(frame-level annotations) (video-level annotations)
(trained on TVSum)
(w/o using TVSum data)
KVS DPP sLSTM SM SMRS Quasi MBF CVS SG VESD DSN LM MINI-Net LR
CHD
Ours
VT 0.353 0.399 0.411 0.415 0.272 0.336 0.295 0.328 0.423 0.447 0.373 0.559 0.785 0.850 -
0.897
VU 0.441 0.453 0.462 0.467 0.324 0.369 0.357 0.413 0.472 0.493 0.441 0.429 0.566 0.714 -
0.657
GA 0.402 0.457 0.463 0.469 0.331 0.342 0.325 0.379 0.475 0.496 0.428 0.612 0.736 0.819 -
0.925
MS 0.417 0.462 0.477 0.478 0.362 0.375 0.412 0.398 0.489 0.503 0.436 0.540 0.753 0.786 -
0.583
PK 0.382 0.437 0.448 0.445 0.289 0.324 0.318 0.354 0.456 0.478 0.411 0.604 0.769 0.802 -
0.847
PR 0.403 0.446 0.461 0.458 0.276 0.301 0.334 0.381 0.473 0.485 0.417 0.475 0.633 0.755 -
0.830
FM 0.397 0.442 0.452 0.451 0.302 0.318 0.365 0.365 0.464 0.487 0.412 0.432 0.612 0.716 -
0.697
Bk 0.342 0.395 0.406 0.407 0.297 0.295 0.313 0.326 0.417 0.441 0.368 0.663 0.756 0.773 -
0.833
BT 0.419 0.464 0.471 0.473 0.314 0.327 0.365 0.402 0.483 0.492 0.435 0.691 0.756 0.786 -
0.875
DS 0.394 0.449 0.455 0.453 0.295 0.309 0.357 0.378 0.466 0.488 0.416 0.626 0.656 0.681 -
0.890
Average 0.398 0.447 0.451 0.461 0.306 0.329 0.345 0.372 0.462 0.481 0.424 0.563 0.702 0.768
0.528
0.783
@@ IV-B Comparisons
Method
YouTube Highlights
TVSum
Ours w/o RASL
0.560
0.616
Ours w/o SA
0.638
0.711
Ours w/o auxiliary
0.644
0.724
Ours w/o vision
0.625
0.614
Ours w/o audio
0.630
0.704
Ours
0.651 0.783
@@
Method YouTube Highlights
TVSum
Summation 0.630
0.700
Concatenation 0.628
0.716
Submodule MLP [10] 0.589
0.604
low-rank fusion [9] 0.621
0.736
@@ IV-C Ablation study
Visual target Audio target YouTube Highlights
TVSum
Pixel mel-spectrogram 0.639
0.603
Pixel Audio FVS [57] 0.648
0.608
Visual FVS [14] mel-spectrogram 0.649
0.700
@@
Method
Training method
Parameters
FLOPs
GIFS
Supervised
440.31M
823.44G
sLSTM
Supervised
2.69M
0.07G
SG
Weakly supervised
18.78M
9.16G
Ours
Unsupervised
8.53 M
0.82G

### v3

# Unsupervised Modality-Transferable Video Highlight Detection with Representation Activation Sequence Learning

@@ Abstract
unsupervised, video highlight detection, multimodal, representation activation sequence.
@@ III-E The Auxiliary Task of Masked FVS Reconstruction
GIFs LSVM RARE LM MINI-Net LR CHD Ours
dog 0.308 0.60 0.49 0.579 0.577 0.554 0.606 0.642
gymnast. 0.335 0.41 0.35 0.417 0.574 0.623 0.711 0.758
parkour 0.540 0.61 0.50 0.670 0.698 0.701 0.742 0.709
skating 0.554 0.62 0.25 0.578 0.522 0.691 0.498 0.456
skiing 0.328 0.36 0.22 0.486 0.539 0.601 0.682 0.665
surfing 0.541 0.51 0.49 0.651 0.593 0.598 0.685 0.667
Average 0.464 0.54 0.38 0.564 0.584 0.630 0.654 0.651
@@ IV-A Evaluation datasets
(frame-level annotations) (video-level annotations) (trained on TVSum) (w/o using TVSum data)
KVS DPP sLSTM SM SMRS Quasi MBF CVS SG VESD DSN LM MINI-Net LR CHD Ours
VT 0.353 0.399 0.411 0.415 0.272 0.336 0.295 0.328 0.423 0.447 0.373 0.559 0.785 0.850 - 0.897
VU 0.441 0.453 0.462 0.467 0.324 0.369 0.357 0.413 0.472 0.493 0.441 0.429 0.566 0.714 - 0.657
GA 0.402 0.457 0.463 0.469 0.331 0.342 0.325 0.379 0.475 0.496 0.428 0.612 0.736 0.819 - 0.925
MS 0.417 0.462 0.477 0.478 0.362 0.375 0.412 0.398 0.489 0.503 0.436 0.540 0.753 0.786 - 0.583
PK 0.382 0.437 0.448 0.445 0.289 0.324 0.318 0.354 0.456 0.478 0.411 0.604 0.769 0.802 - 0.847
PR 0.403 0.446 0.461 0.458 0.276 0.301 0.334 0.381 0.473 0.485 0.417 0.475 0.633 0.755 - 0.830
FM 0.397 0.442 0.452 0.451 0.302 0.318 0.365 0.365 0.464 0.487 0.412 0.432 0.612 0.716 - 0.697
Bk 0.342 0.395 0.406 0.407 0.297 0.295 0.313 0.326 0.417 0.441 0.368 0.663 0.756 0.773 - 0.833
BT 0.419 0.464 0.471 0.473 0.314 0.327 0.365 0.402 0.483 0.492 0.435 0.691 0.756 0.786 - 0.875
DS 0.394 0.449 0.455 0.453 0.295 0.309 0.357 0.378 0.466 0.488 0.416 0.626 0.656 0.681 - 0.890
Average 0.398 0.447 0.451 0.461 0.306 0.329 0.345 0.372 0.462 0.481 0.424 0.563 0.702 0.768 0.528 0.783
@@ IV-B Comparisons
Method YouTube Highlights TVSum
Ours w/o RASL 0.560 0.616
Ours w/o SA 0.638 0.711
Ours w/o auxiliary 0.644 0.724
Ours w/o vision 0.625 0.614
Ours w/o audio 0.630 0.704
Ours 0.651 0.783
@@
Method YouTube Highlights TVSum
Summation 0.630 0.700
Concatenation 0.628 0.716
Submodule MLP [10] 0.589 0.604
low-rank fusion [9] 0.621 0.736
@@ IV-C Ablation study
Visual target Audio target YouTube Highlights TVSum
Pixel mel-spectrogram 0.639 0.603
Pixel Audio FVS [57] 0.648 0.608
Visual FVS [14] mel-spectrogram 0.649 0.700
@@
Method Training method Parameters FLOPs
GIFS Supervised 440.31M 823.44G
sLSTM Supervised 2.69M 0.07G
SG Weakly supervised 18.78M 9.16G
Ours Unsupervised 8.53 M 0.82G